OpenStudy (anonymous):

what's an equation for position vs time graph and a velocity vs time graph using just variables?

OpenStudy (anonymous):

pleasee help!

OpenStudy (anonymous):

position X=Vo t + 0.5at^2 velocity V= Vo + at V^2 = Vo^2 +2as

OpenStudy (anonymous):

where x=s =position, vo = initial velicity v = final velocity a = acceleration t time in seconds

OpenStudy (anonymous):

 position vs time

OpenStudy (anonymous):

 velocity vs time (in sec)

OpenStudy (anonymous):

will that help you out a bit?

OpenStudy (anonymous):

velocity could be accelerating or increasing with time or just decelerating or decreasing with time

OpenStudy (anonymous):

Yes! A lot thanks!